# Formal Verification of Noisy Quantum Reinforcement Learning Policies

### v1

@@ 6. Discussion
QVerifier addresses the high cost and limited availability of quantum hardware [maring2023one, DBLP:journals/sqj/MoguelRVBGM22]: verifying QRL policies classically avoids the expense of on-device testing [DBLP:journals/sqj/MoguelRVBGM22]. QRL policy verification reveals potential safety violations, shows how a specific quantum noise model affects policy behavior, and supports deployment decisions.
A key feature is the exact analytical treatment of quantum noise. Instead of relying on empirical sampling, we compute action probabilities directly from the density matrix, yielding noise-dependent exact transition probabilities. This enables continuous noise-parameter sweeps and precise analysis of how safety properties change with different noise likelihoods (see Figure 6).
The resulting noise–safety profiles guide hardware selection [phalak2025qualiti]. If, for example, a policy remains safe up to a depolarizing rate of $p=0.05$, practitioners can choose devices whose gate fidelities meet this requirement and allocate error budgets accordingly.
A fundamental tension underlies any classical verification of quantum systems [babbush2025grand, meyer2022survey, bowles2024better]: if a circuit is small enough to compute efficiently, one might question the benefit of quantum execution; if it is too large, classical verification becomes too costly [DBLP:journals/corr/abs-2410-12660].
Nevertheless, QVerifier targets a sweet spot: trained QRL policies that run efficiently on quantum hardware but would be far too slow for real-time execution on classical devices, while still remaining computable on classical hardware for model checking, where extended computation time is acceptable. However, this sweet spot remains to be demonstrated in practice; characterizing its boundaries and showing that relevant problems fall within it is essential for validating QVerifier’s real-world applicability.
When full environment verification is infeasible, QVerifier still supports targeted debugging. Practitioners can analyze critical regions of the state space to locate unsafe behaviors.
Another advantage of QVerifier is that it can also be used to analyze different QRL policy designs [meyer2022survey], enabling systematic comparisons [bowles2024better] of how architectural choice, such as circuit depth [fosel2021quantum], encoding schemes [rath2024quantum], or measurement strategies [li2001quantum], impact robustness under noise [DBLP:journals/corr/abs-2410-21117] and adherence to safety requirements, including for smaller quantum circuits. Note that QVerifier can also be used without quantum noise, which is particularly relevant for quantum-inspired reinforcement learning [habibi2025quantum] approaches that employ quantum formalizations in their classical policy design without requiring quantum hardware.
Beyond its immediate applications, QVerifier may inform the design of future quantum model checkers [baltazar2008quantum].
@@ 7. Conclusion
Future work includes integrating explainable RL methods to attribute safety performance to specific gates and circuit elements [DBLP:conf/esann/GrossS24], using model-checking feedback to guide training so that QRL agents learn better policies [gu2024review], and quantum noise mitigation [muqeet2025quiet] for QRL.

### v2

@@ 6. Discussion
QVerifier addresses the high cost and limited availability of quantum hardware [49, 54]: verifying QRL policies classically avoids the expense of on-device testing [54]. QRL policy verification reveals potential safety violations, shows how a specific quantum noise model affects policy behavior, and supports deployment decisions.
A key feature is the exact analytical treatment of quantum noise. Instead of relying on empirical sampling, we compute action probabilities directly from the density matrix, yielding noise-dependent exact transition probabilities. This enables continuous noise-parameter sweeps and precise analysis of how safety properties change with varying noise likelihoods (see Figure 6). When the policy was trained on a simulator with matching noise conditions, the resulting formal model is exact. When trained on physical hardware, however, the model constitutes an idealized approximation, as unknown hardware noise prevents exact modeling. Nevertheless, even idealized models provide valuable insights: they reveal how policies respond to specific noise types and magnitudes, enabling comparative analysis and identification of fragile designs.
The resulting noise–safety profiles guide hardware selection [61]. If, for example, a policy remains safe up to a depolarizing rate of $p=0.05$, practitioners can choose devices whose gate fidelities meet this requirement and allocate error budgets accordingly.
When full environment verification is infeasible [23], QVerifier still supports targeted debugging. Practitioners can analyze critical regions of the state space to locate unsafe behaviors.
A fundamental tension underlies any classical verification of quantum systems [6, 52, 11]: if a circuit is small enough to compute efficiently, one might question the benefit of quantum execution; if it is too large, classical verification becomes too costly [16]. QVerifier targets a sweet spot: trained QRL policies that run efficiently on quantum hardware and are still computable on classical hardware for model checking. This enables QVerifier to analyze different QRL policy designs [52], allowing systematic comparisons [11] of how architectural choices, such as circuit depth [21], encoding schemes [63], or measurement strategies [45], impact robustness under noise [51] and adherence to safety requirements, including for smaller quantum circuits. Note that QVerifier can also be used without quantum noise, which is particularly relevant for quantum-inspired reinforcement learning [30] approaches that employ quantum formalizations in their classical policy design without requiring quantum hardware.
Beyond its immediate applications, QVerifier may inform the design of future quantum model checkers [8].
@@ 7. Conclusion
Future work includes integrating explainable RL methods to attribute safety performance to specific gates and circuit elements [27], using model-checking feedback to guide training so that QRL agents learn better policies [28], and quantum noise mitigation [55] for QRL.